OpenStudy (mirandaabieber):

What is the volume of a sphere with a diameter of 8 meters? Use 3.14 for pi. Round your answer to the nearest hundredth. 33.59 m3 113.10 m3 267.95 m3 2144.66 m3

OpenStudy (3mar):

May I help?

OpenStudy (mirandaabieber):

yes please

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Do you know the sphere volume?

OpenStudy (mirandaabieber):

no. I'm really bad with measurements and volumes...

OpenStudy (3mar):

You will be excellent In Sha' Allah Take that and tell me that you got it.

OpenStudy (mirandaabieber):

I'm so confused

OpenStudy (3mar):

Where are you stuck?

OpenStudy (mirandaabieber):

everywhere if I'm being honest.

OpenStudy (3mar):

Ok Step by step

OpenStudy (3mar):

The volume of the sphere is calculated using the formula \[\frac{ 4 }{ 3 }\pi*r^3\] where r is the radius of the sphere! Got that?

OpenStudy (mirandaabieber):

ok.

OpenStudy (3mar):

Nice! so now we need to substitute with r=8 in that equation and get the volume directly!

OpenStudy (mirandaabieber):

okay i kinda understand

OpenStudy (3mar):

very good! Can I see your substitution!?

OpenStudy (mirandaabieber):

so it becomes ? \[\frac{ 3 }{ 4} \pi \times 8 \]

OpenStudy (3mar):

8 raised to the power 3

OpenStudy (mirandaabieber):

oh okay.

OpenStudy (3mar):

and it is\[\frac{ 4 }{ 3 }\] not \[\frac{ 3 }{ 4 }\]

OpenStudy (mirandaabieber):

ohh so it would be \[\frac{ 4 }{ 3 }\pi \times 8^{3} ?\]

OpenStudy (3mar):

Congratulations! you deserve a medal!

OpenStudy (3mar):

So what is the final result?

OpenStudy (mirandaabieber):

okay so i guess thats where i'm getting stuck i'm not sure how to solve it.

OpenStudy (3mar):

just multiply them together!

OpenStudy (mirandaabieber):

ok give me 1 second.

OpenStudy (3mar):

take it and 10 more

OpenStudy (mirandaabieber):

i got 2,144.6144..... i feel that it's wrong :\

OpenStudy (3mar):

WHy that feeling? Did you use pi as 22/7?

OpenStudy (mirandaabieber):

no. I dont think i did

OpenStudy (3mar):

yes, this result for pi=22/7 and the correct one is 2143.57 m^3 if you use pi=3.14 as he asked. Notice: don't forget the units!!

OpenStudy (mirandaabieber):

oh okay. Thank you so much!!!! :D

OpenStudy (3mar):

You are welcome! Did you get what you are looking for?

OpenStudy (mirandaabieber):

yes i did. thank you for the help i understand this much better now :)

OpenStudy (3mar):

Thanks for Allah. I am happy to hear that Thank you for learning!!

OpenStudy (3mar):

Any more questions?

OpenStudy (mirandaabieber):

nope thats it. :)